# Diffusion Models for Increasing Accuracy in Olfaction Sensors and Datasets

### v1

@@ Abstract
Robotic odour source localization (OSL) is a critical capability for autonomous systems operating in complex environments. However, current OSL methods often suffer from ambiguities, particularly when robots misattribute odours to incorrect objects due to limitations in olfactory datasets and sensor resolutions. To address this challenge, we introduce a novel machine learning method using diffusion-based molecular generation to enhance odour localization accuracy that can be used by itself or with automated olfactory dataset construction pipelines with vision-language models (VLMs) This generative process of our diffusion model expands the chemical space beyond the limitations of both current olfactory datasets and the training data of VLMs, enabling the identification of potential odourant molecules not previously documented. The generated molecules can then be more accurately validated using advanced olfactory sensors which emulate human olfactory recognition through electronic sensor arrays. By integrating visual analysis, language processing, and molecular generation, our framework enhances the ability of olfaction-vision models on robots to accurately associate odours with their correct sources, thereby improving navigation and decision-making in environments where olfactory cues are essential. Our methodology represents a foundational advancement in the field of robotic olfaction, offering a scalable solution to the challenges posed by limited olfactory data and sensor ambiguities.
@@ 1 Introduction
The human sense of smell is a complex and nuanced sensory modality, capable of distinguishing an extensive array of odourants. This capability is underpinned by a sophisticated interplay between olfactory receptors and neural processing pathways. In recent years, artificial olfactory systems, such as electronic noses, have been developed to mimic this biological process, utilizing sensor arrays and pattern recognition algorithms to detect and classify volatile compounds [9].
@@
Despite these advancements, the development of machine learning models for olfactory perception faces significant challenges, primarily due to the scarcity of comprehensive and high-quality olfactory datasets. Datasets like Leffingwell [1] and GoodScents [5] have been instrumental in providing foundational data, comprising thousands of molecules annotated with odour descriptors. However, due to the complexity in obtaining human-labeled odour descriptors over compounds, these datasets are limited in scope, lacking complete coverage of the vast chemical space associated with olfactory stimuli [15].
To address this data paucity, we propose an innovative approach that leverages vision-language models (VLMs) to generate odour descriptors from images. By analyzing visual content in a standard computer vision dataset like, for example, the COCO dataset [12], VLMs can infer potential olfactory characteristics associated with depicted scenes or objects. For example, if a motorcycle is identified in an image, the VLM will reason that carbon monoxide is likely a present molecule and will list the aromas that describe carbon monoxide given the appropriate prompt. These inferred descriptors serve as a bridge to map visual information to chemical compounds, facilitating the identification of molecules likely present in a given image.
However, the reliance on VLMs introduces its own set of uncertainties, particularly regarding the accuracy and completeness of the generated odour descriptors. It is difficult to know the exact scope of chemistry contained in the VLM training data. To mitigate these concerns, we integrate a diffusion model trained on existing olfactory datasets, which is the focus of our work here. This model is designed to generate novel molecular structures that correspond to the inferred odour descriptors, effectively expanding the chemical space beyond the limitations of current datasets. By doing so, we aim to capture a broader spectrum of odourant molecules, including those not previously documented. This is helpful in tasks oriented around olfactory sensors. Sensors that rapidly sample the air only screen for a select set of molecules. If a robot is given the task to navigate toward a specific odour, it has to be equipped with olfactory sensors that can target the specific molecules attributed to that odour. By understanding the target molecule’s near neighbors on the odour continuum, one can ensure to integrate sensors that address these molecules as well. As a simplified example, imagine that an unmanned aerial vehicle (UAV) is instructed to find the source of nitric oxide ($NO$). $NO$ can, in some environments, quickly oxidize to $NO_{2}$. If the time that a robot takes to navigate to the source of $NO$ is substantial, it could be analyzing the wrong compound by missing the detection window, and at some point, may want to transition to detecting $NO_{2}$. The olfactory sensors (for example, metal oxide sensors) required to detect both $NO$ and $NO_{2}$ are different, and it would be prudent for the engineer to query which near neighbor compounds to $NO$ could be relevant for detection so the UAV could be equipped with such sensors. We can make such a query automatic in our VLM dataset construction, such that a separate machine learning model is queried for the nearest neighbors of any compound identified by the VLM. This not only informs better sensor selection, but can also partially reconcile erroneous compounds produced by the VLM.
In machine olfaction, there is a theory that molecules that exhibit similar structural traits "smell" similarly [6] [14] [17] [2]. Our proposed method assumes this theory is true, and we construct a model that builds off its principles. For example, if a VLM indicates that the odour descriptors "fruity" and "floral" are associated with an object in an image, one can use the GoodScent or Leffingwell datasets to look up particular molecules that exude such aromas. However, since these datasets are not comprehensive of all possible molecules, one can use the proposed diffusion model to generate "near neighbors" of molecules that are structurally similar to those found in the existing datasets.
@@
Our research presents a novel methodology that combines visual analysis, language processing, and molecular generation to enhance the field of olfactory machine learning. By addressing the limitations of existing datasets and incorporating advanced sensor validation, we aim to advance the capabilities of artificial olfactory systems in accurately identifying and classifying odourants.
@@ 2 Background
To address the data scarcity issue, researchers have explored the use of vision-language models (VLMs) to generate odour descriptors from images. VLMs, trained on large-scale image-text pairs, can infer contextual information from visual inputs and generate corresponding textual descriptions [19]. By applying VLMs to images, it is possible to extract descriptive terms that may correlate with olfactory characteristics. However, the reliability of these descriptors is contingent upon the VLM’s exposure to olfactory-related data during training, which is often limited. Consequently, the generated descriptors may lack specificity or accuracy in representing actual odours.
@@
Olfaction is a unique sense in that multiple modalities can be used to detect an odourant. Some sensors "see" odourants by measuring the change in optical wavelength bands as a molecule passes through. Other detectors measure small chemical reactions that occur through electron, oxygen, or other molecule transfer across a diffusion medium, such as an electrolyte. Still other sensors "hear" odourants by measuring the vibrational modes of the compound at a quantum level. The first two exemplify what is commonly referred to as the Shape Theory of Olfaction (STO), while the latter alludes to the "Vibrational Theory of Olfaction" (VTO).
First postulated by Dyson in 1928 [6] and again supported in 1938 [14], the vibrational theory of olfaction was originally dismissed as experimental data via Raman spectroscopy was gathered as evidence against the idea. Luca Turin re-popularized the idea in 2001 [17] claiming that Dyson’s proposal was not incorrect, but perhaps slightly uncalibrated. Turin hypothesized that the detection mechanism in mammalian olfactory receptors is due to inelastic electron tunneling. Brooks, et al. discuss the feasibility of a derivative of VTO called phonon-assisted tunneling in [4]. Block, et al. have strongly refuted the plausibility of VTO in [2], leaving an unsettled debate on the matter.
Both VTO and STO provide convincing evidence on why each theory is plausible, and both sides have experimental data to support. "Neither theory can fully explain why the scent of some molecules are concentration dependent. This is a problem well known to perfumers, and yet unexplained" [3]. Stevens reports in his 1951 work that "it is very probable that no one physical property alone is involved in the physical nature of the adequate stimulus" [3]. This leaves the science of olfactory sensing still open with no single sensor variety acting as a unifying chemical sensor.
@@ 3 Related Work
For instance, a study introduced a fusion navigation algorithm that combines both vision and olfaction-based techniques. This hybrid approach addresses challenges such as turbulent airflow, which disrupts olfaction sensing, and physical obstacles inside the search area, which may impede vision detection. The methodology includes a custom-trained deep-learning model for visual target detection and a moth-inspired algorithm for olfaction-based navigation. Experimental results demonstrate that this vision and olfaction fusion algorithm significantly outperforms vision-only and olfaction-only methods, reducing average search time by 54% and 30%, respectively [8] [7].
@@
Furthermore, the application of diffusion models in molecular design has been extensively reviewed, highlighting their potential in generating molecules with desired properties. These models operate by simulating the gradual degradation of a data distribution and learning its reverse process to generate new samples. Their success in visual domains has inspired researchers to explore their potential for molecular generative tasks, making them central to molecular design [18].
@@ 6 Experiments & Results
The first of our experiments was to train the EGNN and diffusion model on the LeffingWell and GoodScent datasets and evaluate the efficacy with the VLM. We borrow the dataset from [11] as it succinctly combines both datasets into a methodical format. The resulting dataset is nearly 5000 samples in size. While this is a small dataset in comparison to modern machine learning tasks in other modalities like image recognition, it is rare to have a dataset this size in the realm of olfaction. We use an 80-20 train-test split in which we evaluate our models over the indicated data. We find that diffusing between 800-1200 steps is ideal, and perform our experiments at 1000 steps. We train the EGNN for 1000 epochs with an embedding dimension of size 8. We train two models in the exact same manner with the exception that one model is unconstrained in that it can generate molecules from any available atom; the other model is "constrained" in that it can only generate molecules from the following atoms: C, N, O, F, P, S, and Cl. The training loss for both models is shown in Figure 3.
@@
We expected to see better performance by constraining the diffusion to a specific set of atoms. However, this actually resulted in worse performance because the permutation of elements did not provide a diverse enough dataset from which the diffusion model could generate neighboring compounds in most cases. While the unconstrained model is more complex, we find that the increased sample space in combination with a longer training time yielded better performance than the constrained model and we leverage that model to communicate our final results.
Although molecules are generated for every set of descriptors that are input to the model, we find that only 27.71% of molecules diffused from our test set pass our molecular validation checks. This is to be expected and quite welcome as we suspect not every permutation of odour descriptors returns a large array of near-neighbor molecules.
@@
Note that our training data does not include all feasible compounds and their odour descriptors  [^1] For more information on the limitations of the Leffingwell and GoodScent datasets, please consult [1] and [5], respectively. We acknowledge that our approach here inherits any limitations noted in these datasets, including those attributed to the subjectivity of human-produced olfactory labels. However, we hope that our diffusion model partially reconciles some of the uncertainty associated with these olfactory labels.. We then test our proposed model with the VLM over a random sample of 1000 images from the COCO dataset. We obtain more permutations of odour descriptors than what are available in our training data, which leads to more possible molecules being produced than those experienced in our training set. Because of this, we note a 28.20% success rate in diffusing near-neighbor molecules from VLM odour descriptors after validation checks are passed.
@@ 7 Limitations
Implementing the proposed framework in real-time robotic systems poses computational challenges. The integration of VLMs, diffusion models, and olfactory sensors requires efficient processing capabilities to ensure timely decision-making during navigation. Latency in processing can hinder the robot’s responsiveness, especially in dynamic environments where rapid adaptation is necessary. It should be noted that the method proposed here is intended to aid in the selection of which olfactory sensors should be integrated on a robot to navigate to a particular compound, prior to deployment on real hardware.
@@ 8 Conclusion
Machine olfaction is still a young area of AI and robotics that receives disproportionate attention and standardization in comparison to other modalities such as computer vision and audition. This creates several opportunities for dataset construction, standardization, and benchmarking. Our methodology represents a foundational step towards improving sensor selection in olfactory robotics by integrating visual, linguistic, and olfactory data. We note opportunities to utilize our method as a standalone algorithm and in an automation pipeline used to construct olfactory datasets with VLMs. It enhances the robot’s ability to generate and associate odour descriptors with potential molecular structures, thereby aiding in more accurate odour localization. However, it is not a comprehensive solution. The limitations outlined above highlight the opportunities for continued research and development to address the complexities inherent in olfactory perception. Future work should focus on expanding olfactory datasets, improving model training for better generalization, standardizing data across various olfaction sensors, and optimizing system integration for real-time applications. This manuscript evaluates one method to reconcile a known shortcoming in generating olfaction-vision datasets, but we hope it inspires more researchers in the field to optimize our solution and address other paucities in machine olfaction.

### v2

# Diffusion Graph Neural Networks for Robustness in Olfaction Sensors and Datasets

@@ Abstract
Robotic odour source localization (OSL) is a critical capability for autonomous systems operating in complex environments. However, current OSL methods often suffer from ambiguities, particularly when robots misattribute odours to incorrect objects due to limitations in olfactory datasets and sensor resolutions. To address this challenge, we introduce a novel machine learning method using diffusion-based molecular generation to enhance odour localization accuracy that can be used by itself or with automated olfactory dataset construction pipelines with vision-language models (VLMs) This generative process of our diffusion model expands the chemical space beyond the limitations of both current olfactory datasets and the training data of VLMs, enabling the identification of potential odourant molecules not previously documented. The generated molecules can then be more accurately validated using advanced olfactory sensors which emulate human olfactory recognition through electronic sensor arrays. By integrating visual analysis, language processing, and molecular generation, our framework enhances the ability of olfaction-vision models on robots to accurately associate odours with their correct sources, thereby improving navigation and decision-making through better sensor selection for a target compound. Our methodology represents a foundational advancement in the field of artificial olfaction, offering a scalable solution to the challenges posed by limited olfactory data and sensor ambiguities.
@@ 1 Introduction
The human sense of smell is a complex and nuanced sensory modality, capable of distinguishing an extensive array of odourants. This capability is underpinned by a sophisticated interplay between olfactory receptors and neural processing pathways. In recent years, artificial olfactory systems, such as electronic noses, have been developed to mimic this biological process, utilizing sensor arrays and pattern recognition algorithms to detect and classify volatile compounds [6, 8, 13, 16].
@@
Despite sensor advancements, the development of machine learning models for olfactory perception faces significant challenges, primarily due to the scarcity of comprehensive olfactory datasets. Datasets like those from Leffingwell [2] and GoodScents [7] have been instrumental in providing foundational results, comprising thousands of molecules annotated with odour descriptors. However, due to the complexity in obtaining human-labeled odour descriptors over compounds, these datasets are limited in scope, lacking complete coverage of the vast chemical space associated with olfactory stimuli [26] and the necessary human controls to ensure objective measurements such as genetic variation in olfactory mechanisms [34], environmental context [23], health status [5, 19], age [10], and even potential neurocognitive conditions [9].
To address this data paucity, we propose an innovative diffusion-based graph neural network that allows one to minimize uncertainty the construction of olfactory datasets, particularly olfaction-vision datasets. Our research provides three primary contributions:
  1. 1.
    Our model aids in the selection of olfactory sensors for a desired target compound.
  2. 2.
    Our model increases the robustness of constructing olfactory-vision datasets through the use of vision-language models (VLMs) to generate odour descriptors from images.
  3. 3.
    In result of (1) and (2), our model reduces overall uncertainty and bias in olfactory tasks.
@@
We combine visual analysis, language processing, and molecular generation to enhance the field of olfactory machine learning through the construction of robust olfaction-vision datasets. By addressing the limitations of existing datasets and incorporating advanced sensor validation, we aim to advance the capabilities of artificial olfactory systems in accurately identifying and classifying odourants.
2 Background
While efforts to construct olfaction-vision-language models (OVLMs) exist [12, 27], their underlying training data and training processes are not well standardized. By analyzing visual content in a standard computer vision dataset like, for example, the COCO dataset [20], VLMs can infer potential olfactory characteristics associated with depicted scenes or objects. For example, if a motorcycle is identified in an image, the VLM will reason that carbon monoxide is likely a present molecule and will list the aromas that describe carbon monoxide given the appropriate prompt. These inferred descriptors serve as a bridge to map visual information to chemical compounds, facilitating the identification of molecules likely present in a given image.
However, the reliance on VLMs introduces its own set of uncertainties, particularly regarding the accuracy and completeness of the generated odour descriptors. It is difficult to know the exact scope of chemistry contained in the VLM training data. To mitigate these concerns, we integrate a diffusion model trained on existing olfactory datasets and designed to generate novel molecular structures that correspond to the inferred odour descriptors, effectively expanding the chemical space beyond the limitations of current datasets. By doing so, we aim to capture a broader spectrum of odourant molecules, including those not previously documented. This is helpful in tasks oriented around olfactory sensors. Sensors that rapidly sample the air only screen for a select set of molecules. If a robot is given the task to navigate toward a specific odour, it has to be equipped with olfactory sensors that can target the specific molecules attributed to that odour. By understanding the target molecule’s near neighbors on the odour continuum, one can ensure to integrate sensors that address these molecules as well. As a simplified example, imagine that an unmanned aerial vehicle (UAV) is instructed to find the source of nitric oxide ($NO$). $NO$ can, in some environments, quickly oxidize to $NO_{2}$. If the time that a robot takes to navigate to the source of $NO$ is substantial, it could be analyzing the wrong compound by missing the detection window, and at some point, may want to transition to detecting $NO_{2}$. The olfactory sensors (for example, metal oxide sensors) required to detect both $NO$ and $NO_{2}$ are different, and it would be prudent for the engineer to query which near neighbor compounds to $NO$ could be relevant for detection so the UAV could be equipped with such sensors. We can make such a query automatic in our VLM dataset construction, such that a separate machine learning model is queried for the nearest neighbors of any compound identified by the VLM. This not only informs better sensor selection, but can also partially reconcile erroneous compounds produced by the VLM.
In machine olfaction, the Shape Theory of Olfaction suggests that molecules that exhibit similar structural traits "smell" similarly [3] [28] [33] . Although there is some support against this theory [4] [11] [24] [30], our proposed method assumes this theory is true, and we construct a model that builds off its principles. For example, if a VLM indicates that the odour descriptors "fruity" and "floral" are associated with an object in an image, one can use the GoodScent or Leffingwell datasets to look up particular molecules that exude such aromas. However, since these datasets are not comprehensive of all possible molecules, one can use the proposed diffusion model to generate "near neighbors" of molecules that are structurally similar to those found in the existing datasets.
@@ 2 Background
To address the data scarcity issue, researchers have explored the use of vision-language models (VLMs) to generate odour descriptors from images. VLMs, trained on large-scale image-text pairs, can infer contextual information from visual inputs and generate corresponding textual descriptions [25] [32]. By applying VLMs to images, it is possible to extract descriptive terms that may correlate with olfactory characteristics. However, the reliability of these descriptors is contingent upon the VLM’s exposure to olfactory-related data during training, the degree of which is not known. Consequently, the generated descriptors may lack specificity or accuracy in representing actual odours.
@@ 3 Related Work
For instance, a study introduced a fusion navigation algorithm that combines both vision and olfaction-based techniques. This hybrid approach addresses challenges such as turbulent airflow, which disrupts olfaction sensing, and physical obstacles inside the search area, which may impede vision detection. The methodology includes a custom-trained deep-learning model for visual target detection and a moth-inspired algorithm for olfaction-based navigation. Experimental results demonstrate that this vision and olfaction fusion algorithm significantly outperforms vision-only and olfaction-only methods, reducing average search time by 30% and 54%, respectively [14] [15].
@@
Furthermore, the application of diffusion models in molecular design has been extensively reviewed, highlighting their potential in generating molecules with desired properties. These models operate by simulating the gradual degradation of a data distribution and learning its reverse process to generate new samples. Their success in visual domains has inspired researchers to explore their potential for molecular generative tasks, making them central to molecular design [31]. Work from Lee, et al. in [18] show how graph neural networks can be used to construct principal odor maps for deducing probable aromas for a given molecule. Our work here with graph neural networks strives to align with many of their assumptions, but leverages diffusion to overcome uncertainties associated with changing perceivable aromas among different concentrations of the same compound as exemplified by Longin, et al. [21].
@@ 6 Experiments & Results
The first of our experiments was to train the EGNN and diffusion model on the LeffingWell and GoodScent datasets and evaluate the efficacy with the VLM. We borrow the dataset from [18] as it succinctly combines both datasets into a methodical format. The resulting dataset is nearly 5000 samples in size. While this is a small dataset in comparison to modern machine learning tasks in other modalities like image recognition, it is rare to have a dataset this size in the realm of olfaction. We use an 80-20 train-test split in which we evaluate our models over the indicated data. We find that diffusing between 800-1200 steps is ideal, and perform our experiments at 1000 steps. We train the EGNN for 1000 epochs with an embedding dimension of size 8. We train two models in the exact same manner with the exception that one model is unconstrained in that it can generate molecules from any available atom; the other model is "constrained" in that it can only generate molecules from the following atoms: C, N, O, F, P, S, and Cl. The training loss for both models is shown in Figure 3. The final results are summarized in Table 1.
@@
We expected to see better performance by constraining the diffusion to a specific set of atoms. However, this actually resulted in worse performance, generating validated molecules over less than 10% of samples. We suspect this is because the permutation of such a small set of elements did not provide a diverse enough dataset from which the diffusion model could generate neighboring compounds in most cases. While the unconstrained model is more complex, we find that the increased sample space in combination with a longer training time yielded better performance than the constrained model and we leverage that model to communicate our final results.
@@
Although molecules are generated for every set of descriptors that are input to the model, we find that only 27.71% of molecules diffused from our test set pass our molecular validation checks. This is to be expected and quite welcome as we suspect not every permutation of identified odour descriptors returns a large array of near-neighbor molecules. Note that our training data does not include all feasible compounds and their odour descriptors  [^1] For more information on the limitations of the Leffingwell and GoodScent datasets, please consult [2] and [7], respectively. We acknowledge that our approach here inherits any limitations noted in these datasets, including those attributed to the subjectivity of human-produced olfactory labels. However, we hope that our diffusion model partially reconciles some of the uncertainty associated with these olfactory labels..
We then test our proposed model with the VLM over the full test data of the COCO dataset [20]. We select GPT-4o [1] as our VLM and prompt it to caption each image with odour descriptors that are suspected to be present. An array of odour descriptors is then associated with each image in the test set and input into our diffusion-EGNN model. We obtain more permutations of odour descriptors than what are available in our training data, which leads to more possible molecules being produced than those experienced in our training set. Because of this, we note a 28.20% success rate in diffusing near-neighbor molecules from VLM odour descriptors after validation checks are passed. This marks a pleasing transfer of training accuracy to test accuracy and amounts evidence for us to suggest that the aromas for the sample of molecules provided in the consolidated Leffingwell and GoodScent datasets may be good proxies of the much larger molecular distribution.
Table 1: Final results: Percentage of test samples with chemically-viable generated molecules.
Dataset Constrained Model Unconstrained Model
GoodScent & LeffingWell Test Set [18] <10% 27.71%
COCO Test Set [20] <10% 28.20%
@@ 7 Limitations
Implementing the proposed framework in real-time robotic systems poses computational challenges. The integration of VLMs, diffusion models, and olfactory sensors requires efficient processing capabilities to ensure timely decision-making during navigation. Latency in processing can hinder the robot’s responsiveness, especially in dynamic environments where rapid adaptation is necessary. It should be noted that the proposed method is intended to aid in the selection of which olfactory sensors should be integrated on a robot to navigate to a particular compound, prior to deployment.
@@ 8 Conclusion
Machine olfaction is still a young area of AI and robotics that receives disproportionate attention and standardization in comparison to other modalities such as computer vision and audition. This creates several opportunities for dataset construction, standardization, and benchmarking. Our methodology represents a foundational step towards improving sensor selection in olfactory robotics by integrating visual, linguistic, and olfactory data. We note opportunities to utilize our method as both a standalone algorithm and as part of an automation pipeline used to construct olfactory datasets via VLMs. Our method enhances the robot’s ability to generate and associate odour descriptors with near-neighbor molecular structures, thereby aiding in more accurate odour localization. However, it is not a comprehensive solution. The limitations outlined above highlight the opportunities for continued research to address the complexities inherent in olfactory perception. This manuscript evaluates one method to reconcile known shortcomings in both the construction of olfaction-vision datasets and the selection of sensors for a target compound, but we hope it inspires more researchers in the field to optimize our solution and address the vast array of opportunities within artificial olfaction.